A standard NaOH solution was prepared by dissolving 0.05 mol of NaOH in 250 mL of

distilled water. 25 mL of the prepared NaOH solution was titrated with 0.15 M H2SO4

solution. Calculate the volume of H2SO4 solution that is required to reach the equivalence

point.


1
Expert's answer
2021-06-25T05:02:03-0400

"H_2SO_4+ 2NaOH \to Na_2SO_4 +H_2O"

molarity of NaOH = "\dfrac{0.05}{250}× 1000 = 0.2M"


"\dfrac{c_Av_A}{c_Bv_B} = \dfrac{n_A}{n_B}"


"\dfrac{0.15× v_A}{0.2×25} =\dfrac12"


"v_A = \dfrac{0.2× 25}{0.15× 2} =\dfrac{5}{0.3} = 16.67mL"